OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (-6, 0)? A.y= (x- 6)2 B.y= (x+ 6)2 C.y=x2- 6 D.y=x2+ 6

OpenStudy (anonymous):

y=a(x-h)^2+k is the equation of parabola in which (h,k) is the vertex

OpenStudy (precal):



OpenStudy (precal):

that vertex is a double root

OpenStudy (anonymous):

so, we have write y= (x+ 6)^2 as y=1(x+6)^2+0

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

now on comparing y=a(x-h)^2+k y=1(x+6)^2+0 we have h=-6 and k=0 so, vertex=(-6,0) for y=(x+6)^2 B) is the correct answer

OpenStudy (anonymous):

yes:)